OpenStudy (anonymous):

FIND 2 numbers that are greater than 0.6 and less than 3/4 the first number (x) must be a fraction and the second number (y) must be a decimal. which number is larger explain why what number is exactly halfway between 0.6 and 3/4.explain how you determined this

OpenStudy (anonymous):

@sweetsunray

OpenStudy (anonymous):

@Satan 

OpenStudy (satan ):

In decimals, 3/4 is 0.75. So look for numbers between 0.60 and 0.75.

OpenStudy (satan ):

To find the exact number in between them, make them fractions, \(0.6=\frac{3}{5}\) and as you know \(0.75=\frac{3}{4}\) and then add them, and multiply by \(\frac{1}{2}\)

OpenStudy (anonymous):

were you answering the second question where i asked is exactly half way between 0.6

OpenStudy (satan ):

Yes, do you understand what to do or do you need me to walk you through it?

OpenStudy (anonymous):

i need you to walk me through finding the 2 numbers greater than 0.6 and less than 3/4

OpenStudy (satan ):

Well look at the numbers: \(\Large 0.6\) and \(\Large 0.75\), what numbers are between it?

OpenStudy (satan ):

Think of it in terms of money if it helps. What is more than 60 cents but less than 75 cents?

OpenStudy (satan ):

Is 50 cents more than 60 cents?

OpenStudy (anonymous):

know um a 1.00

OpenStudy (satan ):

That's bigger than 75 cents :P

OpenStudy (satan ):

Here, I'll help ya out: In terms of 5's: Between .60 and .75 are .65, and 0.7

OpenStudy (anonymous):

so the answer was 65

OpenStudy (satan ):

So your two numbers can be 0.65 and 0.7. You can change 0.7 into a fraction easily, since you need one as a fraction: So use x as \(\Large \frac{7}{10}\) and y as 0.65 for your decimal.

OpenStudy (anonymous):

ohhhh i get it now thanks fro your help

OpenStudy (satan ):

Anytime :)